OpenStudy (anonymous):

Find the greatest common factor of these two expressions. 20w^5v^2y^6 and 24v^8y^4

OpenStudy (anonymous):

@alerion99 can you help me?

OpenStudy (anonymous):

@ticklemepickle. @SweetPeaGirl

OpenStudy (anonymous):

1) Find the greatest common factor of the two expression: Simplify 20uy^8z^6 and 28u^7y^4 --- gcf for 20 and 28 is 4 gcf for u and u^7 is u gcf for y^8 and y^4 is y^4 gcf for z^6 and no z is no z --------------------------------- GCF = 4uy^4 ======================================================================== 2) Find the least common multiple of the two expressions: Simplify 3t^6x^8y^5 and 12x^3y^6 lcm of 3 and 12 is 12 lcm of t^6 and no t is t^6 lcm of x^8 and x^3 is x^8 lcm of y^5 and y^6 is y^6 --- LCM = 12t^6*x^8*y^6 ======================================================== 3) Factor Completely: 27z^3+125 No common factor ------------------------------- 4) Factor: 9t+wt+9y+yw --- t(9+w) + y(9+w) --- = (9+w)(t+y) ------------------------------- 5) Factor: -8t+xt-24x+3x^2 --- = -(8t -xt + 24x - 3x^2) --- = -(t(8-x) + 3x(8-x)) --- = -(8-x)(t+3x)

OpenStudy (anonymous):

holy cow I got lost after two seconds there

jagr2713 (jagr2713):

Hey @alerion99 its better to guide the person to the answer instead of doing to work for them :D

OpenStudy (anonymous):

@jagr2713 do you know where the t came from?

OpenStudy (anonymous):

so what was t originally?

OpenStudy (anonymous):

The t is just a variable I used , dont let it confuse you.

OpenStudy (anonymous):

What grade is this?

OpenStudy (anonymous):

math college readiness 11th or 12th